OpenStudy (anonymous):

Find each percent 34 of 5

OpenStudy (anonymous):

lol troll

OpenStudy (anonymous):

can you copy and paste the question

Directrix (directrix):

@danielcvalencia Percent means hundredths so 34% can be written as 34/100 34% of 5 = 34/100 * 5/1 = 34/20 = 34/20 = 17/10 = 1.7 Question for you: What is 5% of 34?